OpenStudy (anonymous):

Suppose that the sales at Borders bookstores went from 62 million dollars in 1992 to 427 million dollars in 1994. Find an exponential function to model the sales (in millions of dollars) as a function of years (t) since 1992. sales=_______ millions of dollars What is the continuous percent growth rate, per year, of sales? continuous growth rate =

OpenStudy (anonymous):

Have you started it or tried anything yet?

OpenStudy (anonymous):

I dont even know where to begin

OpenStudy (anonymous):

start time at 1992 ie t=0 there...

OpenStudy (anonymous):

so t=2 in 1994

OpenStudy (anonymous):

ok that makes sense

OpenStudy (anonymous):

your expression the looks like 427= 62*e^(c*2)

OpenStudy (anonymous):

so how do i find e and c relative to sales in millions of dollars

OpenStudy (anonymous):

and the general expression is 62*e^(c*t) (you'll need to use the first expression I gave to find 'c'

OpenStudy (anonymous):

hmm so should i put t as 0?

OpenStudy (anonymous):

t=0 in 1992

OpenStudy (anonymous):

i mean to find c

OpenStudy (anonymous):

427= 62*e^(c*2) use natural log to find c

OpenStudy (anonymous):

(and algebra)

OpenStudy (anonymous):

ok sorry my prof touched very lightly on this stuff. So can you tell me what i take the logs of is it log of both sides?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

simplify a bit first to make your life easier...

OpenStudy (anonymous):

so is it log of 427 and log of 62? or is itlog62e^elog2

OpenStudy (anonymous):

\[\ln (e ^{value}) = value\]

OpenStudy (anonymous):

ie natural log 'undoes' e^

OpenStudy (anonymous):

divide both sides by 427 first, then take the natural log of both sides

OpenStudy (anonymous):

so is it 427/427= (ln62*e^(c*2))/427

OpenStudy (anonymous):

wait you said the log gets rid of the e^

OpenStudy (anonymous):

yes... actually I should have said divide both sides by 62 got it turned around... do that and then take the natural log of both sides...

OpenStudy (anonymous):

so then (427/62)=6.8870=....

OpenStudy (anonymous):

=ln2/62?

OpenStudy (anonymous):

.... try using the draw feature to show me what you're doing and I can check it

OpenStudy (anonymous):

ok u said divide both sides by 62 and the first equation you told me to set up was 427= 62*e^(c*2). So i did (427/62)=((ln2)/62) because you said the natural log gets rid of the e^ and you said divide both sides by 62.

OpenStudy (anonymous):